A MNC want to convert (sell) EUR2,000,000 and USD5,000,000 to GBP. Market rates currently are:
EUR1: GBP 0.8680 – 0.8695
USD1: GBP 0.7692 - 0.7705

How much GBP will that MNC receive
Answer #1

A MNC want to convert (sell) EUR2,000,000 and USD5,000,000 to GBP. Market rates currently are:
EUR1: GBP 0.8680 – 0.8695
USD1: GBP 0.7692 - 0.7705

Here number of euros = 20,00,000 EUR

Now the selling rate of Euro to GBP is 0.8680

Number of GBP received = 20,00,000 * 0.8680 = 17,36,000 GBP

Now number of US Dollars = 50,00,000

Now the selling rate of Dollars to GBP is 0.7692

Number of GBP received = 50,00,000 * 0.7692 = 38,46,000 GBP

Total GBP MNC will receive = 1736000 + 3846000 = 55,82,000 GBP

GBP that will MNC receive is 55,82,000 GBP .